OpenStudy (anonymous):

At what value of parameter "a" are there values of x such that the numbers \[5^{1+x}+5^{1-x},\frac{a}{2},25^{x}+25^{-x}\] form an Arithmetic Progression for real x.

OpenStudy (anonymous):

If this a, b and c are in AP then: \[b = \frac{a + c}{2}\]

OpenStudy (anonymous):

Plug in the values there...

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Take the first step don't go for second before the first.. Just plug in the values there and show me..

OpenStudy (anonymous):

\[\frac{a}{2}= \frac{5^{1+x}+5^{1-x}+25^{x}+25^{-x}}{2}\]

OpenStudy (anonymous):

Multiply by 2 both the sides..

OpenStudy (anonymous):

\[a= {5^{1+x}+5^{1-x}+25^{x}+25^{-x}}\]

OpenStudy (anonymous):

You can write 25 as: \[25 = 5^2\]

OpenStudy (anonymous):

\[25^x = 5^{2x}\]

OpenStudy (anonymous):

\[{a}= {5^{1+x}+5^{1-x}+5^{2x}+5^{-2x}}\]

OpenStudy (anonymous):

Yes great.. And you can also write: \[5^{1+x} = 5 \cdot 5^x\]

OpenStudy (anonymous):

\[{a}= {5.5^{x}+5.5^{-x}+5^{2x}+5^{-2x}}\]

OpenStudy (anonymous):

Can you take \(5^x\) common there??

OpenStudy (anonymous):

??

OpenStudy (anonymous):

*Factor out..

OpenStudy (anonymous):

\[{a}= 5^x({5+5.5^{-1}+5^{2}+5^{-2}})\]

OpenStudy (anonymous):

\[{a}= 5^x({5+1+25+\frac{1}{25}}) \implies a = 5^x(31 + \frac{1}{25})\]

OpenStudy (anonymous):

776/25

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

\[a = 5^x(\frac{776}{5^2}) \implies a = 5^{x-2} \times 776\]

OpenStudy (anonymous):

then.

OpenStudy (anonymous):

a belongs to [12,inf) how??

OpenStudy (anonymous):

Wait..

OpenStudy (anonymous):

No my brain is not working.. Don't worry.. @mukushla can you help here..??

OpenStudy (anonymous):

\[t=5^x>0\] then\[a=5(t+\frac{1}{t})+t^2+\frac{1}{t^2}\]

OpenStudy (anonymous):

complete the square or take derivative.

OpenStudy (anonymous):

i dont know derivatives.

OpenStudy (anonymous):

pls explain using completing the sq.

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok i understood

OpenStudy (anonymous):

continue.......

OpenStudy (anonymous):

\[t+\frac{1}{t}=(\sqrt{t}-\frac{1}{\sqrt{t}})^2+2 \ge2\]\[t^2+\frac{1}{t^2}=(t-\frac{1}{t})^2+2 \ge2\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[a\ge 5(2)+2=12\]

OpenStudy (anonymous):

Wow @mukushla ..

OpenStudy (anonymous):

:)